Tarzan wishes to use a light vine to swing from the ground onto a tree branch 3.5m high. How fast should he be going when he grabs the vine?
1
Expert's answer
2020-12-10T11:03:55-0500

According to the energy conservation, Tarzan's initial kinetic energy should be equal to his final potential energy:


"K = U\\
\dfrac{mv^2}{2} = mgh"

where "m" is his mass, "v" is his speed, "g = 9.8m\/s^2" is the gravitational acceleration, and "h = 3.5m" is his final height. Thus, expressing initial velocity, obtain:


"v = \sqrt{2gh}\\
v = \sqrt{2\cdot9.8\cdot 3.5 } \approx 8.3m/s"

Answer. 8.3 m/s.